OpenStudy (anonymous):

Find the eighth term of (3a+b)^9 Homeschooled so any help would be appreciated thanks.

OpenStudy (highschoolmom2010):

this is just 

OpenStudy (anonymous):

easier with the binomial theorem

OpenStudy (whpalmer4):

yeah, the 8th term of that... suppose the problem was to find the 97th term of \[ (3a+b)^{126}\]? :-) the answer is to use the binomial theorem...

OpenStudy (whpalmer4):

the binomial theorem allows you to calculate any term directly, without having to expand that monster (which by the way is:\[6561 a^8+17496 a^7 b+20412 a^6 b^2+13608 a^5 b^3+5670 a^4 b^4\]\[+1512 a^3 b^5+252 a^2 b^6+XXX+b^8\]

OpenStudy (whpalmer4):

(I've blocked out the 8th term so that we can have the fun of finding it)

OpenStudy (whpalmer4):

do you know anything about the binomial theorem?

OpenStudy (anonymous):

OHHH so it is kinda like using pascal's triangle? I haven't learned much about it yet but I will look into it, thanks for the help!

OpenStudy (whpalmer4):

yes, pascal's triangle gives you the binomial coefficients...there's also a way to compute them directly if you don't want to build the triangle.

OpenStudy (anonymous):

I'm a bit confused how wouuld I start?

OpenStudy (whpalmer4):

once you know the binomial coefficient for the term of interest, you just multiply it by the appropriate powers of x and y. say we are doing \[(x+y)^n\](in our problem, \(x = 3a,~ y = b,~n = 8\)) the binomial coefficient is written \(\binom{n}{k}\) and can be found by \[\binom{n}{k} = \frac{n!}{(n-k)!k!}\] where \(n! = n*(n-1)*(n-2)...2*1\) \(k\) is a variable that tells us which term we are finding. It goes from 0 to \(n\). The "appropriate powers of \(x\) and \(y\)" are \[x^{n-k}y^{k}\] So, let's do an easy example: \[(x+2)^3\]We could multiply this out by hand easily enough: \[(x+2)^3 = (x+2)(x+2)(x+2) = (x^4+4x+4)(x+2) = x^3+6x^2+12x+8\] Let's see how we would find the 3rd term with the binomial theorem. \[n = 3, k = 3-1=2, x = x, y = 2\] \[\binom{3}{2} = \frac{3!}{(3-2)!2!} = \frac{3*2*1}{1*2*1} = 3\] Our term is \[\binom{n}{k}x^{n-k}y^{k} = \binom{3}{2}x^{3-2}y^{2} = 3x^1y^2 \]But remember that \(y = 2\) so that becomes \[3x^1(2)^2 = 12x\]Sure enough, \(12x\) is the 3rd term when we consult our hand-multiplied version.

OpenStudy (anonymous):

Thank you! Just got the answer!

OpenStudy (whpalmer4):

In your problem, we have \[x = 3a,~y=b, n = 8, k=8-1 =7\] So the 8th term is \[\binom{8}{7} x^7y^1 = \frac{8!}{(8-7)!7!} (3a)^1(b)^7 = 8*(3a)b^7 = 24ab^7 \]